# Complementary Search of Fermionic Absorption Operators at Hadron Collider and Direct Detection Experiments

### v2

@@ 4 Absorption at Nuclear Targets
$\overline{|\mathcal{M}_{A}|^{2}}\equiv\frac{1}{2s_{\chi}+1}\frac{1}{2J+1}\sum_{\begin{subarray}{c}\text{ initial \& }\\ \text{ final spins }\end{subarray}}|\mathcal{M}_{A}|^{2}\,,$ (4.4)
@@
where $q^{2}\equiv(p_{f}-p_{i})^{2}=(p_{\nu}-p_{\chi})^{2}$ is the squared momentum transfer and $F_{\rm Res}^{N}(q^{2})$ are the corresponding response functions. The amplitude for the DM scattering off a point-like nucleus (PLN) $\mathcal{M}_{\rm PLN}^{N}(q^{2})$ is related to the scattering amplitude $\mathcal{M}_{N}(q^{2})$ off a single nucleon (with nucleon mass normalized to $m_{A}$). The difference between the SI and SD interactions are distinguished by a coherence factor $C_{N}$. For the SI scattering, $C_{N}=Z$ and $A-Z$ for proton and neutron, respectively, while $C_{N}=1$ for its SD counterpart.
@@ 4.2 Spin-Dependent Absorption
[…]
3 keV
Borexino [136] $\mathrm{C}_{6}\mathrm{H}_{3}\left(\mathrm{CH}_{3}\right)_{3}$ 958.58 t×$\times$×yr
[…]
@@
CRESST-III ($\mathrm{Li}\mathrm{Al}\mathrm{O}_{2}$) [144] $\mathrm{Li}\mathrm{Al}\mathrm{O}_{2}$ 2.345 kg×$\times$×day
@@
PICO-60 ($\mathrm{C}_{3}\mathrm{F}_{8}$) [137] $\mathrm{C}_{3}\mathrm{F}_{8}$ 2571 kg×$\times$×day
@@ 4.3 Absorption for the Tensor Operator
Since the nuclear target is unpolarized and the momentum transfer orientation (or equivalently the DM velocity) is isotropic, this interference term would not leave any effect in the total event rate. So the total scattering rate is just incohenrent sum of the SD and SI contributions,
@@ 5 Conclusion
Starting from the quark-level fermionic absorption operators, we study for the first time their sensitivities at the LHC and compare with the constraints from the DM direct detection as well as neutrino experiments. In addition to covering the light DM scenarios, the collider search applies to dark sector fermions in general with much larger mass range extending up to the TeV scale. Even for DM with typical mass below $\mathcal{O}(100)$ MeV, the LHC searches can provide better sensitivities although the identification still requires direct detection experiments that has unique recoil energy spectrum. While for scalar and vector operators, the DM direct detection experiments with heavy nuclei target has much better sensitivities, the remaining pseudo-scalar, axial-vector, and tensor operators has much better sensitivities at neutrino experiments with light nuclei. It is also interesting to observe that the tensor operator can have comparable contributions from both the SI and SD scattering.
@@ Acknowledgements
SFG and KM would like to thank Xiao-Dong Ma for useful discussions. The authors are supported by the National Natural Science Foundation of China (12375101, 12425506, 12090060, 12090064, and 12305110) and the SJTU Double First Class start-up fund (WF220442604). SFG is also an affiliate member of Kavli IPMU, University of Tokyo. KM is supported by the Natural Science Basic Research Program of Shaanxi (Program No. 2023-JC-YB-041).
@@ A.1 Scalar Current
$\sigma_{u}^{p}=(17\pm 5)\,\mathrm{MeV}\,,$ $\sigma_{d}^{p}=(32\pm 10)\,\mathrm{MeV}\,,$ (A.2)
$\sigma_{u}^{n}=(15\pm 5)\,\mathrm{MeV}\,,$ $\sigma_{d}^{n}=(36\pm 10)\,\mathrm{MeV}\,.$
@@ A.5 Tensor Operator
$F_{T,0}^{q/N}\left(q^{2}\right)=m_{q}A_{T,10}^{q/N}\left(q^{2}\right),\quad F_{T,1}^{q/N}\left(q^{2}\right)=-m_{q}B_{T,10}^{q/N}\left(q^{2}\right),\quad F_{T,2}^{q/N}\left(q^{2}\right)=\frac{m_{q}}{2}\tilde{A}_{T,10}^{q/N}\left(q^{2}\right).$ (A.20)
[…]
@@
$-0.2\lesssim B_{T,10}^{s/p}(0),\quad\tilde{A}_{T,10}^{s/p}(0)\lesssim 0.2$ (A.23)
@@
[…]
@@ Appendix B Non-Relativistic Expansion
$\frac{\overline{u_{D}}(\vec{p}_{1},s)}{\sqrt{E+m}}=\left(\begin{array}[]{cc}\xi_{s}^{\dagger}&\xi_{s}^{\dagger}\dfrac{-\vec{\sigma}\cdot\vec{p}_{1}}{E+m}\end{array}\right)\,,\;\;\;\frac{\overline{v_{D}}(\vec{p}_{2},r)}{\sqrt{E+m}}=\left(\begin{array}[]{cc}r\xi_{-r}^{\dagger}\dfrac{\vec{\sigma}\cdot\vec{p}_{2}}{E+m}\quad\quad-r\xi_{-r}^{\dagger}\end{array}\right)\,,$ (B.15)

### v3

@@ 4 Absorption at Nuclear Targets
The above Sec. 3 deals with the DS fermions that have absorption operator in general which is a great advantage of collider searches. The DM particle that survives until today in our Universe is just a part of the DS particle zoo. It is then possible to establish complementary searches between the collider and direct detection experiments. However, the DM particle needs to be stable enough. For a fermionic absorption DM, its mass cannot be too large and is typically below 100 MeV [11, 12, 124, 125, 126]. In the current section, we would focus on this light mass regime.
@@
[…]
@@
where $q^{2}\equiv(p_{f}-p_{i})^{2}=(p_{\nu}-p_{\chi})^{2}$ is the squared momentum transfer and $F_{\rm Res}^{N}(q^{2})$ are the corresponding response functions. The amplitude for the DM scattering off a point-like nucleus (PLN) $\mathcal{M}_{\rm PLN}^{N}(q^{2})$ is related to the scattering amplitude $\mathcal{M}_{N}(q^{2})$ off a single nucleon (with nucleon mass normalized to $m_{A}$). The difference between the SI and SD interactions are distinguished by a coherence factor $C_{N}$. For the SI scattering, $C_{N}=Z$ and AZA-Z for proton and neutron, respectively, while $C_{N}=1$ for its SD counterpart.
@@ 4.2 Spin-Dependent Absorption
C  $^{13}_{\;\,6}\mathrm${C} (1.1%)
@@
[…]
@@ 4.3 Absorption for the Tensor Operator
Since the nuclear target is unpolarized and the momentum transfer orientation (or equivalently the DM velocity) is isotropic, this interference term would not leave any effect in the total event rate.
So the total scattering rate is just incohenrent sum of the SD and SI contributions,
@@ 5 Conclusion
Starting from the quark-level fermionic absorption operators, we study for the first time their sensitivities at the LHC and compare with the constraints from the DM direct detection as well as neutrino experiments. While the LHC can probe the DS particles in general, the collider and direct detection experiments can provide complementary searches for the DM particle that survives until today with small enough mass. The sensitivity at LHC can reach $\Lambda\gtrsim\mathcal{O}(1)$ TeV with mono-$X$ (photon, jet, and $Z$) channels. Even for DM with typical mass below $\mathcal{O}(100)$ MeV, the LHC searches can provide better sensitivities although the identification still requires direct detection experiments that has unique recoil energy spectrum. While for scalar and vector operators, the DM direct detection experiments with heavy nuclei target has much better sensitivities, the remaining pseudo-scalar, axial-vector, and tensor operators has much better sensitivities at neutrino experiments with light nuclei. It is also interesting to observe that the tensor operator can have comparable contributions from both the SI and SD scattering.
@@ Acknowledgements
SFG and KM would like to thank Xiao-Dong Ma for useful discussions. The authors are supported by the National Natural Science Foundation of China (12375101, 12425506, 12090060, 12090064, and 12305110) and the SJTU Double First Class start-up fund (WF220442604). SFG is also an affiliate member of Kavli IPMU, University of Tokyo.
@@ A.1 Scalar Current
[…]
@@ A.5 Tensor Operator
$\hskip-5.69054ptF_{T,0}^{q/N}\left(q^{2}\right)=m_{q}A_{T,10}^{q/N}\left(q^{2}\right),\quad F_{T,1}^{q/N}\left(q^{2}\right)=-m_{q}B_{T,10}^{q/N}\left(q^{2}\right),\quad F_{T,2}^{q/N}\left(q^{2}\right)=\frac{m_{q}}{2}\tilde{A}_{T,10}^{q/N}\left(q^{2}\right).$ (A.20)
At the leading order, only $F_{T,0}^{q/N}(0)$ and $F_{T,1}^{q/N}(0)$ appear. The value of $F_{T,0}^{q/N}(0)$ is quite well determined and is usually expressed in terms of $A_{T,10}^{q/p}(0)\equiv g_{T}^{q}$ (with $A_{T,10}^{u(d)/p}=$ $A_{T,10}^{d(u)/n}$ and $A_{T,10}^{s/p}=A_{T,10}^{s/n}$ in the isospin limit). The tensor charges $g_{T}^{q}$ are related to the transversity structure functions $\delta q_{N}(x,\mu)$ by $g_{T}^{q}(\mu)=$ $\int_{-1}^{1}dx\delta q_{N}(x,\mu)$. The lattice calculations in the $\overline{\mathrm{MS}}$ scheme at $\mu=2\,\mathrm{GeV}$, including both connected and disconnected contributions gives [168, 169],
@@
$0.2\lesssim B_{T,10}^{s/p}(0),\quad\tilde{A}_{T,10}^{s/p}(0)\lesssim 0.2$ (A.23)
@@
$\sum_{q=u,d,s}\frac{F_{T,0}^{q/p}(0)}{m_{q}}\approx 0.59,\quad\sum_{q=u,d,s}\frac{F_{T,1}^{q/p}(0)}{m_{q}}\in\big[-2.96,\,-2.56\big]\,,\quad\sum_{q=u,d,s}\frac{F_{T,2}^{q/p}(0)}{m_{q}}\in\big[-0.12,\,0.08\big]\,$,
@@ Appendix B Non-Relativistic Expansion
$\frac{\overline{u_{D}}(\vec{p}_{1},s)}{\sqrt{E+m}}=\left(\begin{array}[]{cc}\xi_{s}^{\dagger}\quad\quad&\xi_{s}^{\dagger}\dfrac{-\vec{\sigma}\cdot\vec{p}_{1}}{E+m}\end{array}\right)\,,\;\;\;\frac{\overline{v_{D}}(\vec{p}_{2},r)}{\sqrt{E+m}}=\left(\begin{array}[]{cc}r\xi_{-r}^{\dagger}\dfrac{\vec{\sigma}\cdot\vec{p}_{2}}{E+m}\quad\quad-r\xi_{-r}^{\dagger}\end{array}\right)\,$, (B.15)